OpenStudy (anonymous):

Solve the equation x/6 + 9 = 14

OpenStudy (anonymous):

x/6+9=14 x/6=14-9 x/6=5 x=30

OpenStudy (callisto):

shouldn't you try do it yourself after we've answered similar questions from you?